# Groupoid Graded Semisimple Rings

### v1

@@ 5.3. Structure of gr-simple $\Gamma_{0}$-artinian rings
It is well known that if $e$ is an idempotent of a ring $X$ and $M$ is a right $X$-module, then $\operatorname{Hom}_{R}(eR,M)\rightarrow Me$, $f\mapsto f(e)$, is an isomorphism of additive groups and $\operatorname{Hom}_{R}(eR,eR)\rightarrow eRe$, $f\mapsto f(e)$, is an isomorphism of rings. We apply this to the idempotent $E^{r(\tau_{j})}_{jj}$ of the ring $R$.
@@
Let $D$ be a gr-division ring and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ a matricial sequence for $D$. Consider the $\Gamma$-graded ring $R:=\operatorname{M}_{I}(D)(\overline{\Sigma})$. The following assertions hold true.
@@
    :=$\mathcal{F}:=\{E_{ii}^{r(\sigma_{i})}R:i\in I,\sigma_{i}\in\Sigma_{i}\}$ is a family of gr-simple $\Gamma$-graded $R$-modules. Analogously, :=$\mathcal{F}^{\prime}:=\{RE_{ii}^{r(\sigma_{i})}:i\in I,\sigma_{i}\in\Sigma_{i}\}$ is a family of gr-simple left $R$-modules.
@@
The following assertions hold true.
@@
:=$\mathcal{F}:=\{T\textrm{ graded right ideal of $R$}:\textrm{there exists $\sigma\in e\Gamma$ such that $T_{R}\cong_{gr}S(\sigma)$}\}.$
@@
Note that, in Proposition 5.20, $D$ is graded by the isotropy group $e\Gamma e$ and we have $r(\sigma_{i})=e$ for all $i\in I$. But the $\Gamma$-grading on $\operatorname{M}_{I}(D)(\overline{\sigma})$ may not be realized as a group grading (in the sense of [11, Definition 1.8]). In fact, if there exist $i\neq j$ such that $d(\sigma_{i})\neq d(\sigma_{j})$, then $E^{e}_{ii}\in\operatorname{M}_{I}(D)(\overline{\sigma})_{d(\sigma_{i})}$ and $E^{e}_{jj}\in\operatorname{M}_{I}(D)(\overline{\sigma})_{d(\sigma_{j})}$. Therefore, $E^{e}_{ii}$ and $E^{e}_{jj}$ are in distinct homogeneous components. This cannot occur in a group grading because $E^{e}_{ii}$ and $E^{e}_{jj}$ are nonzero homogeneous idempotents and thus both should be in the component corresponding to the unity of the group.
On the other hand, the $\Gamma$-grading on $\operatorname{M}_{I}(D)(\overline{\sigma})$ is a refinement of a group grading, see [9, Corollary 2 of Section 4]. Indeed, for each $f\in\Gamma^{\prime}_{0}:=\Gamma^{\prime}_{0}(\operatorname{M}_{I}(D)(\overline{\sigma}))$, take $i_{f}\in I$ such that $d(\sigma_{i_{f}})=f$. Then $\operatorname{M}_{I}(D)(\overline{\sigma})$ is a $e\Gamma e$-graded ring if we define the component of degree $g\in e\Gamma e$ as $\bigoplus\limits_{f,f^{\prime}\in\Gamma^{\prime}_{0}}\operatorname{M}_{I}(D)(\overline{\sigma})_{\sigma_{i_{f}}^{-1}g\sigma_{i_{f^{\prime}}}}$. More generally, suppose that $R=\bigoplus\limits_{\gamma\in\Gamma}R_{\gamma}$ is a $\Gamma$-graded ring such that there is $e\in\Gamma_{0}^{\prime}(R)$ with the property that for all $f\in\Gamma^{\prime}_{0}(R)$ there exists $\sigma_{f}\in e\Gamma f$. In other words, the groupoid $\Gamma^{\prime}:=\{\gamma\in\Gamma:1_{r(\gamma)},1_{d(\gamma)}\neq 0\}$ is connected. Then $R$ is a $e\Gamma e$-graded ring if we define the component of degree $g\in e\Gamma e$ as $\bigoplus\limits_{f,f^{\prime}\in\Gamma^{\prime}_{0}(R)}R_{\sigma_{f}^{-1}g\sigma_{f^{\prime}}}$. This is a coarsening of the $\Gamma$-grading of $R$.
@@
Inspired by [17, Subsection 1.4.1], the next example shows that there does not exist a full version of Proposition 5.24 for the subrings $R_{e}$, $e\in\Gamma_{0}$. In Proposition 5.26, we give a characterization of when the subrings $R_{e}$ are simple artinian.
@@
Let $I_{1},...,I_{n}$ be the equivalence classes of this relation. For each $k=1,...,n$ and $i,j\in I_{k}$, we have, by Proposition 3.5(2),
@@
and it follows that $D(\sigma_{j})\cong_{gr}D(\sigma_{i})$. For each $k=1,...,n$, fix $i_{k}\in I_{k}$ and let $n_{k}:=|I_{k}|$. Set $\overline{\sigma^{\prime}}:=(\sigma^{\prime}_{i})_{i\in I}\in(e_{0}\Gamma)^{I}$ where $\sigma^{\prime}_{i}=\sigma_{i_{k}}$ if $i\in I_{k}$ and $\sigma_{i}^{\prime}=\sigma_{i}$ if $i\notin I_{e}$. Then
@@ 5.4. Structure of gr-semisimple rings
(2) Follows from (1).
@@
Before obtaining some important cases where the converse of items (1)–(3) in Proposition 5.32 hold, we need some definitions. Let $e\in\Gamma_{0}$. For a $\Gamma$-graded ring $R$, we say that $R$ is right $e$-faithful if, for each, $\gamma\in e\Gamma$ and $0\neq a\in R_{\gamma}$, there exists $r\in R_{\gamma^{-1}}$ such that $0\neq ar\in R_{e}$. We will say that $R(e)$ is $\{e\}$-faithful if, for each $0\neq a\in\operatorname{h}(R(e))$, there exists $r\in\operatorname{h}(R)$ such that $0\neq ar\in 1_{e}R1_{e}$. Clearly, if $R$ is right $e$-faithful, then $R(e)$ is $\{e\}$-faithful. We also say that $R$ is strongly $\Gamma$-graded if $R_{\gamma}R_{\delta}=R_{\gamma\delta}$ for all $\gamma,\delta\in\Gamma$. It is not difficult to see that if $R$ is strongly $\Gamma$-graded, then $R$ is right $e$-faithful for all $e\in\Gamma_{0}$. The previous concepts generalize items (2) and (3) of [3, Definition 7 (p. 536)].
@@
(1) Suppose that $e\in\Gamma_{0}$, $R$ is right $e$-faithful and $1_{e}=s_{1}+\cdots+s_{n}$, where $s_{1},...,s_{n}\in R_{e}$ and $s_{i}R_{e}$ is a simple $R_{e}$-module for each $i=1,...,n$. We will show that each $s_{i}R$ is a gr-simple $R$-module. Fix $i=1,...,n$ and $0\neq a\in s_{i}R$. By $e$-faithfulness, there exists $r\in\operatorname{h}(R)$ such that $0\neq ar\in R_{e}$. Since $0\neq ar\in(s_{i}R)_{e}=s_{i}R_{e}$, it follows from simplicity that $arR_{e}=s_{i}R_{e}$. In particular, $s_{i}\in arR_{e}\subseteq aR$ and thus $aR=s_{i}R$. Hence $R(e)=1_{e}R=\sum_{i=1}^{n}s_{i}R$ is a gr-semisimple $R$-module.
(2) Suppose that $e\in\Gamma_{0}$, $R(e)$ is $\{e\}$-faithful and $1_{e}=s_{1}+\cdots+s_{n}$, where $s_{1},...,s_{n}\in R_{e}$ and $s_{i}(1_{e}R1_{e})$ is a gr-simple $1_{e}R1_{e}$-module for all $i=1,...,n$. It suffices to show that each $s_{i}R$ is a gr-simple $R$-module. Fix $i=1,...,n$ and $0\neq a\in s_{i}R$. Take $r\in\operatorname{h}(R)$ such that $0\neq ar\in 1_{e}R1_{e}$. Then $0\neq ar\in 1_{e}(s_{i}R)1_{e}=s_{i}(1_{e}R1_{e})$ and it follows that $ar(1_{e}R1_{e})=s_{i}(1_{e}R1_{e})$. Thus, $s_{i}\in aR$ and therefore $aR=s_{i}R$. ∎
@@
The following assertions are equivalent for a strongly $\Gamma$-graded ring $R$:
@@
The following assertions are equivalent for a gr-prime strongly $\Gamma$-graded ring $R$:
@@
Following [7, Definition 12], we will say that the $\Gamma$-graded ring $R$ is an object crossed product if, for each $\gamma\in\Gamma$, there exists an invertible element in $R_{\gamma}$. Object crossed products are strongly graded. In fact, for each $\sigma,\tau\in\Gamma$ with $d(\sigma)=r(\tau)$, taking an invertible element $u\in R_{\tau}$, we have
@@
Furthermore, [7, Proposition 16], says that the object crossed products are precisely the $\Gamma$-graded rings of the form $A\rtimes^{\alpha}_{\beta}\Gamma$, where $(A,\Gamma,\alpha,\beta)$ is an object crossed system as in Example 2.2(3). In this context, we have:
@@
Let $(A,\Gamma,\alpha,\beta)$ be an object crossed system. The following assertions hold true.
@@
We observe that Corollary 5.38 applies for object twisted groupoid rings [7, Definition 22].
@@
First notice that $\Sigma=\bigcup\limits_{\xi\in\Xi}\Sigma_{\xi}$ and this union is disjoint. Let $(i,\sigma_{i})\in\Sigma$ and take (the unique) $\xi\in\Xi$ such that $r(\sigma_{i})\in\xi$. Thus, $[r(\sigma_{\xi})]=[r(\sigma_{i})]$, i.e., there exists $\gamma_{i,\sigma_{i}}\in(\operatorname{supp}D)\cap r(\sigma_{\xi})\Gamma r(\sigma_{i})$. Then, taking $0\neq u_{i,\sigma_{i}}\in D_{\gamma_{i,\sigma_{i}}}$ we have that
@@
Let $D$ be a $\Gamma$-graded division ring and consider the gr-primality relation $\sim$ defined on $\Gamma^{\prime}_{0}:=\Gamma^{\prime}_{0}(D)$. Then, for each $[e]\in\Gamma^{\prime}_{0}/\!\sim$, there exists $\overline{\gamma}_{[e]}:=(\gamma_{f})_{f\in[e]}\in\prod\limits_{f\in[e]}\operatorname{supp}(1_{e}D1_{f})$ such that
@@
By Remark 3.15(1), $D\cong_{gr}\operatorname{M}_{1}(D)(\Gamma^{\prime}_{0})$. Now apply Proposition 5.39 for $\Xi=\Gamma^{\prime}_{0}/\!\sim$, $\Sigma=\{1\}\times\Gamma^{\prime}_{0}\cong\Gamma^{\prime}_{0}$, $\Sigma_{[e]}=\{1\}\times[e]\cong[e]$ and $\sigma_{[e]}=e$ for each $[e]\in\Gamma^{\prime}_{0}/\!\sim$. ∎
@@
    Any $e\Gamma$-graded right $R$-module is gr-semisimple for some $e\in\Gamma_{0}$.
@@
The equivalence of (1) and (2) follows from Proposition 5.9 and Lemma 3.18, once we observe that a $\Gamma$-graded right $R$-module $M$ is gr-semisimple if and only if $M(e)$ is gr-semisimple for every $e\in\Gamma_{0}$.

### v2

@@ 5.3. Structure of gr-simple $\Gamma_{0}$-artinian rings
It is well-known that if $e$ is an idempotent of a ring $X$ and $M$ is a right $X$-module, then $\operatorname{Hom}_{R}(eR,M)\rightarrow Me$, $f\mapsto f(e)$, is an isomorphism of additive groups and $\operatorname{Hom}_{R}(eR,eR)\rightarrow eRe$, $f\mapsto f(e)$, is an isomorphism of rings. We apply this to the idempotent $E^{r(\tau_{j})}_{jj}$ of the ring $R$.
@@
Let $D$ be a gr-division ring and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ a matricial sequence for $D$. Consider the $\Gamma$-graded ring $R:=\operatorname{M}_{I}(D)(\overline{\Sigma})$. The following assertions hold:
@@
    $\mathcal{F}:=\{E_{ii}^{r(\sigma_{i})}R:i\in I,\sigma_{i}\in\Sigma_{i}\}$ is a family of gr-simple $\Gamma$-graded $R$-modules. Analogously, $\mathcal{F}^{\prime}:=\{RE_{ii}^{r(\sigma_{i})}:i\in I,\sigma_{i}\in\Sigma_{i}\}$ is a family of gr-simple left $R$-modules.
@@
The following assertions hold:
@@
[…]
@@
Note that, in Proposition 5.20, $D$ is graded by the isotropy group $e\Gamma e$ and we have $r(\sigma_{i})=e$ for all $i\in I$. But the $\Gamma$-grading on $\operatorname{M}_{I}(D)(\overline{\sigma})$ may not be realized as a group grading (in the sense of [14, Definition 1.8]). In fact, if there exist $i\neq j$ such that $d(\sigma_{i})\neq d(\sigma_{j})$, then $E^{e}_{ii}\in\operatorname{M}_{I}(D)(\overline{\sigma})_{d(\sigma_{i})}$ and $E^{e}_{jj}\in\operatorname{M}_{I}(D)(\overline{\sigma})_{d(\sigma_{j})}$. Therefore, $E^{e}_{ii}$ and $E^{e}_{jj}$ are in distinct homogeneous components. This cannot occur in a group grading because $E^{e}_{ii}$ and $E^{e}_{jj}$ are nonzero homogeneous idempotents and thus both should be in the component corresponding to the unity of the group.
On the other hand, the $\Gamma$-grading on $\operatorname{M}_{I}(D)(\overline{\sigma})$ is a refinement of a group grading, see [12, Corollary 2 of Section 4]. Indeed, for each $f\in\Gamma^{\prime}_{0}:=\Gamma^{\prime}_{0}(\operatorname{M}_{I}(D)(\overline{\sigma}))$, take $i_{f}\in I$ such that $d(\sigma_{i_{f}})=f$. Then $\operatorname{M}_{I}(D)(\overline{\sigma})$ is a $e\Gamma e$-graded ring if we define the component of degree $g\in e\Gamma e$ as $\bigoplus\limits_{f,f^{\prime}\in\Gamma^{\prime}_{0}}\operatorname{M}_{I}(D)(\overline{\sigma})_{\sigma_{i_{f}}^{-1}g\sigma_{i_{f^{\prime}}}}$. More generally, suppose that $R=\bigoplus\limits_{\gamma\in\Gamma}R_{\gamma}$ is a $\Gamma$-graded ring such that there is $e\in\Gamma_{0}^{\prime}(R)$ with the property that for all $f\in\Gamma^{\prime}_{0}(R)$ there exists $\sigma_{f}\in e\Gamma f$. In other words, the groupoid $\Gamma^{\prime}:=\{\gamma\in\Gamma:1_{r(\gamma)},1_{d(\gamma)}\neq 0\}$ is connected. Then $R$ is a $e\Gamma e$-graded ring if we define the component of degree $g\in e\Gamma e$ as $\bigoplus\limits_{f,f^{\prime}\in\Gamma^{\prime}_{0}(R)}R_{\sigma_{f}^{-1}g\sigma_{f^{\prime}}}$. This is a coarsening of the $\Gamma$-grading of $R$.
@@
Inspired by [20, Subsection 1.4.1], the next example shows that there does not exist a full version of Proposition 5.24 for the subrings $R_{e}$, $e\in\Gamma_{0}$. In Proposition 5.26, we give a characterization of when the subrings $R_{e}$ are simple artinian.
@@
Let $I_{1},...,I_{n}$ be the equivalence classes of this relation. For each k=1,,nk=1,...,n and $i,j\in I_{k}$, we have, by Proposition 3.5(2),
@@
and it follows that $D(\sigma_{j})\cong_{gr}D(\sigma_{i})$. For each k=1,,nk=1,...,n, fix $i_{k}\in I_{k}$ and let $n_{k}:=|I_{k}|$. Set $\overline{\sigma^{\prime}}:=(\sigma^{\prime}_{i})_{i\in I}\in(e_{0}\Gamma)^{I}$ where $\sigma^{\prime}_{i}=\sigma_{i_{k}}$ if $i\in I_{k}$ and $\sigma_{i}^{\prime}=\sigma_{i}$ if $i\notin I_{e}$. Then
@@ 5.4. Structure of gr-semisimple rings
(2) Set $\Delta_{0}=\{e\}$. Since $1_{\Delta_{0}}R1_{\Delta_{0}}=1_{e}R1_{e}$, the result follows from (1).
@@
Before obtaining some important cases where the converse of items (1)–(3) in Proposition 5.32 hold, we need some definitions. Let $e\in\Gamma_{0}$. For a $\Gamma$-graded ring $R$, we say that $R$ is right $e$-faithful if, for each, $\gamma\in e\Gamma$ and $0\neq a\in R_{\gamma}$, there exists $r\in R_{\gamma^{-1}}$ such that $0\neq ar\in R_{e}$. We will say that $R(e)$ is $\{e\}$-faithful if, for each $0\neq a\in\operatorname{h}(R(e))$, there exists $r\in\operatorname{h}(R)$ such that $0\neq ar\in 1_{e}R1_{e}$. Clearly, if $R$ is right $e$-faithful, then $R(e)$ is $\{e\}$-faithful. We also say that $R$ is strongly $\Gamma$-graded if $R_{\gamma}R_{\delta}=R_{\gamma\delta}$ for all $\gamma,\delta\in\Gamma$. It is not difficult to see that if $R$ is strongly $\Gamma$-graded, then $R$ is right $e$-faithful for all $e\in\Gamma_{0}$. The previous concepts generalize items (2) and (3) of [6, Definition 7 (p. 536)].
@@
(1) Suppose that $e\in\Gamma_{0}$, $R$ is right $e$-faithful and $1_{e}=s_{1}+\cdots+s_{n}$, where $s_{1},...,s_{n}\in R_{e}$ and $s_{i}R_{e}$ is a simple $R_{e}$-module for each i=1,,ni=1,...,n. We will show that each $s_{i}R$ is a gr-simple $R$-module. Fix i=1,,ni=1,...,n and $0\neq a\in s_{i}R$. By $e$-faithfulness, there exists $r\in\operatorname{h}(R)$ such that $0\neq ar\in R_{e}$. Since $0\neq ar\in(s_{i}R)_{e}=s_{i}R_{e}$, it follows from simplicity that $arR_{e}=s_{i}R_{e}$. In particular, $s_{i}\in arR_{e}\subseteq aR$ and thus $aR=s_{i}R$. Hence $R(e)=1_{e}R=\sum_{i=1}^{n}s_{i}R$ is a gr-semisimple $R$-module.
(2) Suppose that $e\in\Gamma_{0}$, $R(e)$ is $\{e\}$-faithful and $1_{e}=s_{1}+\cdots+s_{n}$, where $s_{1},...,s_{n}\in R_{e}$ and $s_{i}(1_{e}R1_{e})$ is a gr-simple $1_{e}R1_{e}$-module for all i=1,,ni=1,...,n. It suffices to show that each $s_{i}R$ is a gr-simple $R$-module. Fix i=1,,ni=1,...,n and $0\neq a\in s_{i}R$. Take $r\in\operatorname{h}(R)$ such that $0\neq ar\in 1_{e}R1_{e}$. Then $0\neq ar\in 1_{e}(s_{i}R)1_{e}=s_{i}(1_{e}R1_{e})$ and it follows that $ar(1_{e}R1_{e})=s_{i}(1_{e}R1_{e})$. Thus, $s_{i}\in aR$ and therefore $aR=s_{i}R$. ∎
@@
For a strongly $\Gamma$-graded ring $R$, the following assertions are equivalent:
@@
For a gr-prime strongly $\Gamma$-graded ring $R$, the following assertions are equivalent:
@@
Following [10, Definition 12], we will say that the $\Gamma$-graded ring $R$ is an object crossed product if, for each $\gamma\in\Gamma$, there exists an invertible element in $R_{\gamma}$. Object crossed products are strongly graded. In fact, for each $\sigma,\tau\in\Gamma$ with $d(\sigma)=r(\tau)$, taking an invertible element $u\in R_{\tau}$, we have
@@
Furthermore, [10, Proposition 16], says that the object crossed products are precisely the $\Gamma$-graded rings of the form $A\rtimes^{\alpha}_{\beta}\Gamma$, where $(A,\Gamma,\alpha,\beta)$ is an object crossed system as in Example 2.2(3). In this context, we have:
@@
Let $(A,\Gamma,\alpha,\beta)$ be an object crossed system. The following assertions hold:
@@
We observe that Corollary 5.38 applies for object twisted groupoid rings [10, Definition 22].
@@
First notice that $\Sigma=\bigcup\limits_{\xi\in\Xi}\Sigma_{\xi}$ and this is a union of disjoint sets. Let $(i,\sigma_{i})\in\Sigma$ and take (the unique) $\xi\in\Xi$ such that $r(\sigma_{i})\in\xi$. Thus, $r(\sigma_{\xi})]=[r(\sigma_{i})]$, i.e., there exists $\gamma_{i,\sigma_{i}}\in(\operatorname{supp}D)\cap r(\sigma_{\xi})\Gamma r(\sigma_{i})$. Then, taking $0\neq u_{i,\sigma_{i}}\in D_{\gamma_{i,\sigma_{i}}}$ we have that
@@
Let $D$ be a $\Gamma$-graded division ring and consider the gr-primality relation $\sim$ defined on $\Gamma^{\prime}_{0}:=\Gamma^{\prime}_{0}(D)$. Then, for each $e]\in\Gamma^{\prime}_{0}/\!\sim$, there exists $\overline{\gamma}_{[e]}:=(\gamma_{f})_{f\in[e]}\in\prod\limits_{f\in[e]}\operatorname{supp}(1_{e}D1_{f})$ such that
@@
By Remark 3.16(1), $D\cong_{gr}\operatorname{M}_{1}(D)(\Gamma^{\prime}_{0})$. Now apply Proposition 5.39 for $\Xi=\Gamma^{\prime}_{0}/\!\sim$, $\Sigma=\{1\}\times\Gamma^{\prime}_{0}\cong\Gamma^{\prime}_{0}$, $\Sigma_{[e]}=\{1\}\times[e]\cong[e]$ and $\sigma_{[e]}=e$ for each $e]\in\Gamma^{\prime}_{0}/\!\sim$. ∎
@@
    There exists $e\in\Gamma_{0}$ such that any graded right $R$-module in the full subcategory $e\Gamma-\operatorname{gr\textrm{-}\!}R$ of $\Gamma-\operatorname{gr\textrm{-}\!}R$ is gr-semisimple.
@@
The equivalence of (1) and (2) follows from Proposition 5.9 and Lemma 3.19, upon observing that a $\Gamma$-graded right $R$-module $M$ is gr-semisimple if and only if $M(e)$ is gr-semisimple for every $e\in\Gamma_{0}$.